# Video Prediction Transformers without Recurrence or Convolution

### v3

@@ Abstract
Video prediction has witnessed the emergence of RNN-based models led by ConvLSTM, and CNN-based models led by SimVP. Following the significant success of ViT, recent works have integrated ViT into both RNN and CNN frameworks, achieving improved performance. While we appreciate these prior approaches, we raise a fundamental question: Is there a simpler yet more effective solution that can eliminate the high computational cost of RNNs while addressing the limited receptive fields and poor generalization of CNNs? How far can it go with a simple pure transformer model for video prediction? In this paper, we propose PredFormer, a framework entirely based on Gated Transformers. We provide a comprehensive analysis of 3D Attention in the context of video prediction. Extensive experiments demonstrate that PredFormer delivers state-of-the-art performance across four standard benchmarks. The significant improvements in both accuracy and efficiency highlight the potential of PredFormer as a strong baseline for real-world video prediction applications. The source code and trained models will be released at https://github.com/yyyujintang/PredFormer.
@@ 1 Introduction
Video Prediction [36, 4, 9], also named as Spatio-Temporal predictive learning [34, 35, 27, 28] involves learning spatial and temporal patterns by predicting future frames based on past observations. This capability is essential for various applications, including weather forecasting [23, 22, 3], traffic flow prediction [7, 37], precipitation nowcasting [25, 8] and human motion forecasting [43, 33].
@@
Despite the success of various video prediction methods, they often struggle to balance computational cost and performance. On the one hand, high-powered recurrent-based methods [25, 34, 37, 4, 40, 29, 30] rely heavily on autoregressive RNN frameworks, which face significant limitations in parallelization and computational efficiency. On the other hand, efficient recurrent-free methods [9, 27], such as those based on the SimVP framework, use CNNs in an encoder-decoder architecture but are constrained by local receptive fields, limiting their scalability and generalization. The ensuing question is Can we develop a framework that autonomously learns spatiotemporal dependencies without relying on inductive bias?
@@
An intuitive solution directly adopts a pure transformer [32] structure, as it is an efficient alternative to RNNs and has better scalability than CNNs. Transformers have demonstrated remarkable success in visual tasks [6, 17, 2, 1, 31]. Previous video prediction methods try to combine Swin Transformer [17] in recurrent-based frameworks such as SwinLSTM [29] and integrate MetaFormer [41] as a temporal translator in recurrent-free CNN-based encoder-decoder frameworks such as OpenSTL [28]. Despite these advances, pure transformer-based architecture remains underexplored mainly due to the challenges of capturing spatial and temporal relationships within a unified framework. While merging spatial and temporal dimensions and applying full attention is conceptually straightforward, it is computationally expensive because of the quadratic scaling of attention with sequence length. Several recent methods [2, 1, 31] decouple full attention and show that spatial and temporal relations can be treated separately in a factorized or interleaved manner to reduce complexity.
@@
In this work, we propose PredFormer, a pure transformer-based architecture for video prediction. PredFormer dives into the decomposition of spatial and temporal transformers, integrating self-attention with gated linear units [5] to more effectively capture complex spatiotemporal dynamics. In addition to retaining spatial-temporal full attention encoder and factorized encoder strategies for both spatial-first and temporal-first configurations, we introduce six novel interleaved spatiotemporal transformer architectures, resulting in nine configurations. We explore how far this simple framework can go with different strategies of 3D Attention. This comprehensive investigation pushes the boundaries of current models and sets valuable benchmarks for spatial-temporal modeling.
@@ 2 Related Work
Recurrent-based video prediction. Recent advancements in recurrent-based video prediction models have integrated CNNs, ViTs, and Vision Mamba [16] into RNNs, employing various strategies to capture spatiotemporal relationships. ConvLSTM [25], evolving from FC-LSTM [26], innovatively integrates convolutional operations into the LSTM framework. PredNet [19] leverages deep recurrent convolutional neural networks with bottom-up and top-down connections to predict future video frames. PredRNN [34] introduces the Spatiotemporal LSTM (ST-LSTM) unit, which effectively captures and memorizes spatial and temporal representations by propagating hidden states horizontally and vertically. PredRNN++ [35] incorporates a gradient highway unit and Causal LSTM to address the vanishing gradient problem and adaptively capture temporal dependencies. E3D-LSTM [36] extends the memory capabilities of ST-LSTM by integrating 3D convolutions. The MIM model [37] further refines the ST-LSTM by reimagining the forget gate with dual recurrent units and utilizing differential information between hidden states. CrevNet [40] employs a CNN-based reversible architecture to decode complex spatiotemporal patterns. PredRNNv2 [38] enhances PredRNN by introducing a memory decoupling loss and a curriculum learning strategy. MAU [4] adds a motion-aware unit to capture dynamic motion information. SwinLSTM [29] integrates the Swin Transformer [17] module into the LSTM architecture, while VMRNN [30] extends this by incorporating the Vision Mamba module. Unlike these approaches, PredFormer is a recurrent-free method that offers superior efficiency.
Recurrent-free video prediction. Recent recurrent-free models, e.g., SimVP [9], are developed based on a CNN-based encoder-decoder with a temporal translator. TAU [27] builds upon this by separating temporal attention into static intra-frame and dynamic inter-frame components, introducing a differential divergence loss to supervise inter-frame variations. OpenSTL [28] integrates a MetaFormer model as the temporal translator. Additionally, PhyDNet [10] incorporates physical principles into CNN architectures, while DMVFN [13] introduces a dynamic multi-scale voxel flow network to enhance video prediction performance. EarthFormer [8] presents a 2D CNN encoder-decoder architecture with cuboid attention.WAST [21] proposes a wavelet-based method, coupled with a wavelet-domain High-Frequency Focal Loss. In contrast to prior methods, PredFormer advances video prediction with its recurrent-free, pure transformer-based architecture, leveraging a global receptive field to achieve superior performance, outperforming prior models without relying on complex architecture designs or specialized loss.
@@
Vision Transformer (ViT). ViT [6] has demonstrated exceptional performance on various vision tasks. In video processing, TimeSformer [2] investigates the factorization of spatial and temporal self-attention and proposes that divided attention where temporal and spatial attention are applied separately yields the best accuracy. ViViT [1] explores factorized encoders, self-attention, and dot product mechanisms, concluding that a factorized encoder with spatial attention applied first performs better. On the other hand, TSViT [31] finds that a factorized encoder prioritizing temporal attention achieves superior results. Latte [20] investigates factorized encoders and factorized self-attention mechanisms, incorporating both spatial-first and spatial-temporal parallel designs, within the context of latent diffusion transformers for video generation. Despite these advancements, most existing models focus primarily on video classification, with limited research on applying ViTs to spatio-temporal predictive learning. Moving beyond earlier methods that focus on factorizing self-attention, PredFormer explores the decomposition of spatial and temporal transformers at a deeper level by integrating self-attention with gated linear units and introducing innovative interleaved designs, allowing for a more robust capture of complex spatiotemporal dynamics.
@@ 3.2 Gated Transformer Block
The Standard Transformer model [32] alternates between Multi-Head Attention (MSA) and Feed-Forward Networks (FFN). The attention mechanism for each head is defined as:
@@
Gated Linear Units (GLUs) [5], often used in place of simple linear transformations, involve the element-wise product of two linear projections, with one projection passing through a sigmoid function. Various GLU variants control the flow of information by substituting the sigmoid with other non-linear functions. For instance, SwiGLU [24] replaces the sigmoid with the Swish activation function (SiLU) [11], as shown in Eq 3.2.
SwiGLU has been demonstrated to outperform Multi-layer Perceptrons (MLPs) in various natural language processing tasks[24]. Inspired by the SwiGLU’s success in these tasks, our Gated Transformer Block (GTB), incorporates MSA followed by a SwiGLU-based FFN, as illustrated in Fig 3(a). GTB is defined as:
@@ 3.3 Variants of PredFormer
A PredFormer layer is a module capable of simultaneously processing spatial and temporal information. Building on this design principle, we propose three interleaved spatiotemporal paradigms, Binary, Triplet, and Quadrupled, which sequentially model the relation from spatial and temporal views.
Ultimately, they yield six distinct architectural configurations. A detailed illustration of these nine variants is provided in Fig 3.
@@
For Binary layers, each GTB block processes temporal or spatial sequence independently, which we denote as a binary-TS or binary-ST layer. The input is first reshaped, and processed through , where attention is applied over the temporal sequence. The tensor is then reshaped back to restore the temporal order. Subsequently, spatial attention is applied using another , where the tensor is flattened along the temporal dimension and processed.
For Triplet and Quadruplet layers, additional blocks are stacked on top of the Binary structure. Triplet-TST captures more temporal dependencies, while Triplet-STS focuses more on spatial dependencies, both using the same number of parameters. The Quadruplet layer combines two Binary layers in different orders.
@@ 4 Experiments
Implementation Details Our method is implemented in PyTorch. The experiments are conducted on a single 24GB NVIDIA RTX 3090.
PredFormer is optimized using the AdamW [18] optimizer with an L2 loss, a weight decay of 1e-2, and a learning rate selected from {5e-4, 1e-3} for best performance. OneCycle scheduler is used for Moving MNIST and TaxiBJ, while the Cosine scheduler is applied for Human3.6m and WeatherBench. Dropout [12] and stochastic depth [14] regularization prevent overfitting. Further hyperparameter details are provided in the Appendix. For different PredFormer variants, we maintain a constant number of GTB blocks to ensure comparable parameters. In cases where the Triplet model cannot be evenly divided, we use the number of GTB blocks closest to the others.
Evaluation Metrics We assess model performance using a suite of metrics. (1) Pixel-wise error is measured using Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE). (2) Predicted frame quality is evaluated using the structural similarity index measure (SSIM) metric [39]. Lower MSE, MAE, and RMSE values, combined with higher SSIM, signify better predictions. (3) Computational efficiency is assessed by the number of parameters, floating-point operations (FLOPs), and inference speed in frames per second (FPS) on a NVIDIA A6000 GPU. This evaluation framework comprehensively evaluates accuracy and efficiency.
| Dataset | Training size | Testing size | Channel | Height | Width | Interval | ||
| Moving MNIST | 10,000 | 10,000 | 1 | 64 | 64 | 10 | 10 | - |
| Human3.6m | 73404 | 8582 | 3 | 256 | 256 | 4 | 4 | frame |
| WeatherBench-S | 52559 | 17495 | 1 | 32 | 64 | 12 | 12 | 30 min |
| TaxiBJ | 20,461 | 500 | 2 | 32 | 32 | 4 | 4 | 1 hour |
@@ 4.1 Synthetic Moving Object Prediction
| Method | Paras(M) | Flops(G) | FPS | MSE | MAE | SSIM |
| ConvLSTM | 15.0 | 56.8 | 113 | 103.3 | 182.9 | 0.707 |
| PredRNN | 23.8 | 116.0 | 54 | 56.8 | 126.1 | 0.867 |
| PredRNN++ | 38.6 | 171.7 | 38 | 46.5 | 106.8 | 0.898 |
| MIM | 38.0 | 179.2 | 37 | 44.2 | 101.1 | 0.910 |
| E3D-LSTM | 51.0 | 298.9 | 18 | 41.3 | 86.4 | 0.910 |
| PhyDNet | 3.1 | 15.3 | 182 | 24.4 | 70.3 | 0.947 |
| MAU | 4.5 | 17.8 | 201 | 27.6 | 86.5 | 0.937 |
| PredRNNv2 | 24.6 | 708.0 | 24 | 48.4 | 129.8 | 0.891 |
| SwinLSTM | 20.2 | 69.9 | 62 | 17.7 | - | 0.962 |
| SimVP | 58.0 | 19.4 | 209 | 23.8 | 68.9 | 0.948 |
| TAU | 44.7 | 16.0 | 283 | 19.8 | 60.3 | 0.957 |
| OpenSTL_ViT | 46.1 | 16.9 | 290 | 19.0 | 60.8 | 0.955 |
| OpenSTL_Swin | 46.1 | 16.9 | 290 | 18.3 | 59.0 | 0.960 |
| PredFormer | ||||||
| Full Attention | 25.3 | 21.2 | 254 | 17.3 | 56.0 | 0.962 |
| Fac-S-T | 25.3 | 16.5 | 368 | 20.6 | 63.5 | 0.955 |
| Fac-T-S | 25.3 | 16.5 | 370 | 16.9 | 55.8 | 0.963 |
| Binary-TS | 25.3 | 16.5 | 301 | 12.8 | 46.1 | 0.972 |
| Binary-ST | 25.3 | 16.5 | 316 | 13.4 | 47.1 | 0.971 |
| Triplet-TST | 25.3 | 16.4 | 312 | 13.4 | 47.2 | 0.971 |
| Triplet-STS | 25.3 | 16.5 | 321 | 13.1 | 46.7 | 0.972 |
| Quadruplet-TSST | 25.3 | 16.5 | 302 | 12.4 | 44.6 | 0.973 |
| Quadruplet-STTS | 25.3 | 16.4 | 322 | 12.4 | 44.9 | 0.973 |
Moving MNIST. The moving MNIST dataset [26] serves as a benchmark synthetic dataset for evaluating sequence reconstruction models. We follow [26] to generate Moving MNIST sequences with 20 frames, using the initial 10 frames for input and the subsequent 10 frames as the prediction target. We adopt 10000 sequences for training, and for fair comparisons, we use the pre-generated 10000 sequences [9] for validation.
On the Moving MNIST dataset, following prior work [9, 27], we train our models for 2000 epochs and report our results in Tab 2. We train OpenSTL methods with ViT and Swin Transformer as temporal translators for 2000 epochs as recurrent-free baselines. We cite other results from each original paper for a fair comparison.
@@
Compared to SwinLSTM, PredFormer achieves higher accuracy. Although SwinLSTM outperforms SimVP in terms of MSE (17.7 vs. 23.8), its reliance on an RNN-based structure results in significantly higher computational cost. SwinLSTM exhibits high FLOPs (69.9G) and lower FPS, making it less efficient for large-scale deployment. This highlights the limitations of recurrent structures in video prediction, whereas PredFormer, with its recurrence-free transformer framework, achieves both higher accuracy and superior efficiency.

### v4

@@ Abstract
Video prediction has witnessed the emergence of RNN-based models led by ConvLSTM, and CNN-based models led by SimVP. Following the significant success of ViT, recent works have integrated ViT into both RNN and CNN frameworks, achieving improved performance. While we appreciate these prior approaches, we raise a fundamental question: Is there a simpler yet more effective solution that can eliminate the high computational cost of RNNs while addressing the limited receptive fields and poor generalization of CNNs? How far can it go with a simple pure transformer model for video prediction? In this paper, we propose PredFormer, a framework entirely based on Gated Transformers. We provide a comprehensive analysis of 3D Attention in the context of video prediction. Extensive experiments demonstrate that PredFormer delivers state-of-the-art performance across four standard benchmarks. The significant improvements in both accuracy and efficiency highlight the potential of PredFormer as a strong baseline for real-world video prediction applications. The source code and trained models are released at https://github.com/yyyujintang/PredFormer.
@@ 1 Introduction
Video Prediction (Wang et al., 2018c; Chang et al., 2021; Gao et al., 2022a), also named as Spatio-Temporal predictive learning (Wang et al., 2017; 2018b; Tan et al., 2023a; b) involves learning spatial and temporal patterns by predicting future frames based on past observations. This capability is essential for various applications, including weather forecasting (Rasp et al., 2020; Pathak et al., 2022; Bi et al., 2023), traffic flow prediction (Fang et al., 2019; Wang et al., 2019), precipitation nowcasting (Shi et al., 2015; Gao et al., 2022b) and human motion forecasting (Zhang et al., 2017b; Wang et al., 2018a).
@@
Despite the success of various video prediction methods, they often struggle to balance computational cost and performance. On the one hand, high-powered recurrent-based methods (Shi et al., 2015; Wang et al., 2017; 2019; Chang et al., 2021; Yu et al., 2019; Tang et al., 2023; 2024) rely heavily on autoregressive RNN frameworks, which face significant limitations in parallelization and computational efficiency. On the other hand, efficient recurrent-free methods (Gao et al., 2022a; Tan et al., 2023a), such as those based on the SimVP framework, use CNNs in an encoder-decoder architecture but are constrained by local receptive fields, limiting their scalability and generalization. The ensuing question is Can we develop a framework that autonomously learns spatiotemporal dependencies without relying on inductive bias?
@@
An intuitive solution directly adopts a pure transformer (Vaswani et al., 2017) structure, as it is an efficient alternative to RNNs and has better scalability than CNNs. Transformers have demonstrated remarkable success in visual tasks (Dosovitskiy et al., 2020; Liu et al., 2021; Bertasius et al., 2021; Arnab et al., 2021; Tarasiou et al., 2023). Previous video prediction methods try to combine Swin Transformer (Liu et al., 2021) in recurrent-based frameworks such as SwinLSTM (Tang et al., 2023) and integrate MetaFormer (Yu et al., 2022) as a temporal translator in recurrent-free CNN-based encoder-decoder frameworks such as OpenSTL (Tan et al., 2023b). Despite these advances, pure transformer-based architecture remains underexplored mainly due to the challenges of capturing spatial and temporal relationships within a unified framework. While merging spatial and temporal dimensions and applying full attention is conceptually straightforward, it is computationally expensive because of the quadratic scaling of attention with sequence length. Several recent methods (Bertasius et al., 2021; Arnab et al., 2021; Tarasiou et al., 2023) decouple full attention and show that spatial and temporal relations can be treated separately in a factorized or interleaved manner to reduce complexity.
@@
In this work, we propose PredFormer, a pure transformer-based architecture for video prediction. PredFormer dives into the decomposition of spatial and temporal transformers, integrating self-attention with gated linear units (Dauphin et al., 2017) to more effectively capture complex spatiotemporal dynamics. In addition to retaining spatial-temporal full attention encoder and factorized encoder strategies for both spatial-first and temporal-first configurations, we introduce six novel interleaved spatiotemporal transformer architectures, resulting in nine configurations. We explore how far this simple framework can go with different strategies of 3D Attention. This comprehensive investigation pushes the boundaries of current models and sets valuable benchmarks for spatial-temporal modeling.
@@ 2 Related Work
Recurrent-based video prediction. Recent advancements in recurrent-based video prediction models have integrated CNNs, ViTs, and Vision Mamba (Liu et al., 2024) into RNNs, employing various strategies to capture spatiotemporal relationships. ConvLSTM (Shi et al., 2015), evolving from FC-LSTM (Srivastava et al., 2015), innovatively integrates convolutional operations into the LSTM framework. PredNet (Lotter et al., 2017) leverages deep recurrent convolutional neural networks with bottom-up and top-down connections to predict future video frames. PredRNN (Wang et al., 2017) introduces the Spatiotemporal LSTM (ST-LSTM) unit, which effectively captures and memorizes spatial and temporal representations by propagating hidden states horizontally and vertically. PredRNN++ (Wang et al., 2018b) incorporates a gradient highway unit and Causal LSTM to address the vanishing gradient problem and adaptively capture temporal dependencies. E3D-LSTM (Wang et al., 2018c) extends the memory capabilities of ST-LSTM by integrating 3D convolutions. The MIM model (Wang et al., 2019) further refines the ST-LSTM by reimagining the forget gate with dual recurrent units and utilizing differential information between hidden states. CrevNet (Yu et al., 2019) employs a CNN-based reversible architecture to decode complex spatiotemporal patterns. PredRNNv2 (Wang et al., 2022) enhances PredRNN by introducing a memory decoupling loss and a curriculum learning strategy. MAU (Chang et al., 2021) adds a motion-aware unit to capture dynamic motion information. SwinLSTM (Tang et al., 2023) integrates the Swin Transformer (Liu et al., 2021) module into the LSTM architecture, while VMRNN (Tang et al., 2024) extends this by incorporating the Vision Mamba module. Unlike these approaches, PredFormer is a recurrent-free method that offers superior efficiency.
Recurrent-free video prediction. Recent recurrent-free models, e.g., SimVP (Gao et al., 2022a), are developed based on a CNN-based encoder-decoder with a temporal translator. TAU (Tan et al., 2023a) builds upon this by separating temporal attention into static intra-frame and dynamic inter-frame components, introducing a differential divergence loss to supervise inter-frame variations. OpenSTL (Tan et al., 2023b) integrates a MetaFormer model as the temporal translator. Additionally, PhyDNet (Guen and Thome, 2020) incorporates physical principles into CNN architectures, while DMVFN (Hu et al., 2023) introduces a dynamic multi-scale voxel flow network to enhance video prediction performance. EarthFormer (Gao et al., 2022b) presents a 2D CNN encoder-decoder architecture with cuboid attention.WAST (Nie et al., 2024) proposes a wavelet-based method, coupled with a wavelet-domain High-Frequency Focal Loss. In contrast to prior methods, PredFormer advances video prediction with its recurrent-free, pure transformer-based architecture, leveraging a global receptive field to achieve superior performance, outperforming prior models without relying on complex architecture designs or specialized loss.
@@
Vision Transformer (ViT). ViT (Dosovitskiy et al., 2020) has demonstrated exceptional performance on various vision tasks. In video processing, TimeSformer (Bertasius et al., 2021) investigates the factorization of spatial and temporal self-attention and proposes that divided attention where temporal and spatial attention are applied separately yields the best accuracy. ViViT (Arnab et al., 2021) explores factorized encoders, self-attention, and dot product mechanisms, concluding that a factorized encoder with spatial attention applied first performs better. On the other hand, TSViT (Tarasiou et al., 2023) finds that a factorized encoder prioritizing temporal attention achieves superior results. Latte (Ma et al., 2024) investigates factorized encoders and factorized self-attention mechanisms, incorporating both spatial-first and spatial-temporal parallel designs, within the context of latent diffusion transformers for video generation. Despite these advancements, most existing models focus primarily on video classification, with limited research on applying ViTs to spatio-temporal predictive learning. Moving beyond earlier methods that focus on factorizing self-attention, PredFormer explores the decomposition of spatial and temporal transformers at a deeper level by integrating self-attention with gated linear units and introducing innovative interleaved designs, allowing for a more robust capture of complex spatiotemporal dynamics.
@@ 3.2 Gated Transformer Block
The Standard Transformer model (Vaswani et al., 2017) alternates between Multi-Head Attention (MSA) and Feed-Forward Networks (FFN). The attention mechanism for each head is defined as:
@@
Gated Linear Units (GLUs) (Dauphin et al., 2017), often used in place of simple linear transformations, involve the element-wise product of two linear projections, with one projection passing through a sigmoid function. Various GLU variants control the flow of information by substituting the sigmoid with other non-linear functions. For instance, SwiGLU (Shazeer, 2020) replaces the sigmoid with the Swish activation function (SiLU) (Hendrycks and Gimpel, 2016), as shown in Eq 3.2.
SwiGLU has been demonstrated to outperform Multi-layer Perceptrons (MLPs) in various natural language processing tasks(Shazeer, 2020). Inspired by the SwiGLU’s success in these tasks, our Gated Transformer Block (GTB), incorporates MSA followed by a SwiGLU-based FFN, as illustrated in Fig 3(a). GTB is defined as:
@@ 3.3 Variants of PredFormer
A PredFormer layer is a module capable of simultaneously processing spatial and temporal information. Building on this design principle, we propose three interleaved spatiotemporal paradigms, Binary, Triplet, and Quadruplet, which sequentially model the spatial and temporal views. Ultimately, they yield six distinct architectural configurations. A detailed illustration of these nine variants is provided in Fig 3.
@@
For Binary layers, each GTB block processes temporal or spatial sequence independently, which we denote as a binary-TS or binary-ST layer.
The input is first reshaped, and processed through , where attention is applied over the temporal sequence. The tensor is then reshaped back to restore the temporal order. Subsequently, spatial attention is applied using another , where the tensor is flattened along the temporal dimension and processed.
For Triplet and Quadruplet layers, additional blocks are stacked on top of the Binary structure.
The Quadruplet layer combines two Binary layers in different orders.
@@ 4 Experiments
| Dataset | Training size | Testing size | Channel | Height | Width | Input | Output | Interval |
| Moving MNIST | 10,000 | 10,000 | 1 | 64 | 64 | 10 | 10 | - |
| Human3.6m | 73,404 | 8,582 | 3 | 256 | 256 | 4 | 4 | frame |
| WeatherBench-S | 52,559 | 17,495 | 1 | 32 | 64 | 12 | 12 | 30 min |
| TaxiBJ | 20,461 | 500 | 2 | 32 | 32 | 4 | 4 | 1 hour |
@@
Implementation Details Our method is implemented in PyTorch. The experiments were conducted on a single 24GB NVIDIA RTX 3090. PredFormer is optimized using the AdamW (Loshchilov and Hutter, 2019) optimizer with an L2 loss, a weight decay of 1e-2, and a learning rate selected from {5e-4, 1e-3} for best performance. OneCycle scheduler is used for Moving MNIST and TaxiBJ, while the Cosine scheduler is applied for Human3.6m and WeatherBench. Dropout (Hinton, 2012) and stochastic depth (Huang et al., 2016) regularization prevent overfitting. Further hyperparameter details are provided in the Appendix. For different PredFormer variants, we maintain a constant number of GTB blocks to ensure comparable parameters. In cases where the Triplet model cannot be evenly divided, we use the number of GTB blocks closest to the others.
Evaluation Metrics We assess model performance using a suite of metrics. (1) Pixel-wise error is measured using Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Error (RMSE). (2) Predicted frame quality is evaluated using the structural similarity index measure (SSIM) metric (Wang et al., 2004). Lower MSE, MAE, and RMSE values, combined with higher SSIM, signify better predictions. (3) Computational efficiency is assessed by the number of parameters, floating-point operations (FLOPs), and inference speed in frames per second (FPS) on a NVIDIA A6000 GPU. This evaluation framework comprehensively evaluates accuracy and efficiency.
@@ 4.1 Synthetic Moving Object Prediction
| Method | Paras(M) | Flops(G) | FPS | Memory (MB) | MSE | MAE | SSIM |
| ConvLSTM | 15.0 | 56.8 | 113 | 68.8 | 103.3 | 182.9 | 0.707 |
| PredRNN | 23.8 | 116.0 | 54 | 107.6 | 56.8 | 126.1 | 0.867 |
| PredRNN++ | 38.6 | 171.7 | 38 | 164.1 | 46.5 | 106.8 | 0.898 |
| MIM | 38.0 | 179.2 | 37 | 160.6 | 44.2 | 101.1 | 0.910 |
| E3D-LSTM | 51.0 | 298.9 | 18 | 270.1 | 41.3 | 86.4 | 0.910 |
| PhyDNet | 3.1 | 15.3 | 182 | 149.5 | 24.4 | 70.3 | 0.947 |
| MAU | 4.5 | 17.8 | 201 | 35.4 | 27.6 | 86.5 | 0.937 |
| PredRNNv2 | 24.6 | 708.0 | 24 | 109.0 | 48.4 | 129.8 | 0.891 |
| SwinLSTM | 20.2 | 69.9 | 62 | 96.4 | 17.7 | - | 0.962 |
| SimVP | 58.0 | 19.4 | 209 | 284.7 | 23.8 | 68.9 | 0.948 |
| TAU | 44.7 | 16.0 | 283 | 322.9 | 19.8 | 60.3 | 0.957 |
| OpenSTL_ViT | 46.1 | 16.9 | 290 | 331.8 | 19.0 | 60.8 | 0.955 |
| OpenSTL_Swin | 46.1 | 16.9 | 290 | 331.9 | 18.3 | 59.0 | 0.960 |
| PredFormer | |||||||
| Full Attention | 25.3 | 21.2 | 254 | 135.9 | 17.3 | 56.0 | 0.962 |
| Fac-S-T | 25.3 | 16.5 | 368 | 117.4 | 20.6 | 63.5 | 0.955 |
| Fac-T-S | 25.3 | 16.5 | 370 | 117.4 | 16.9 | 55.8 | 0.963 |
| Binary-TS | 25.3 | 16.5 | 301 | 117.4 | 12.8 | 46.1 | 0.972 |
| Binary-ST | 25.3 | 16.5 | 316 | 117.4 | 13.4 | 47.1 | 0.971 |
| Triplet-TST | 25.3 | 16.4 | 312 | 118.0 | 13.4 | 47.2 | 0.971 |
| Triplet-STS | 25.3 | 16.5 | 321 | 118.0 | 13.1 | 46.7 | 0.972 |
| Quadruplet-TSST | 25.3 | 16.5 | 302 | 118.0 | 12.4 | 44.6 | 0.973 |
| Quadruplet-STTS | 25.3 | 16.4 | 322 | 118.0 | 12.4 | 44.9 | 0.973 |
Moving MNIST. The moving MNIST dataset (Srivastava et al., 2015) serves as a benchmark synthetic dataset for evaluating sequence reconstruction models. We follow (Srivastava et al., 2015) to generate Moving MNIST sequences with 20 frames, using the initial 10 frames for input and the subsequent 10 frames as the prediction target. We adopt 10000 sequences for training, and for fair comparisons, we use the pre-generated 10000 sequences (Gao et al., 2022a) for validation.
On the Moving MNIST dataset, following prior work (Gao et al., 2022a; Tan et al., 2023a), we train our models for 2000 epochs and report our results in Tab 2. We train OpenSTL methods with ViT and Swin Transformer as temporal translators for 2000 epochs as recurrent-free baselines. We cite other results from each original paper for a fair comparison.
@@
Compared to SwinLSTM, PredFormer achieves higher accuracy. Although SwinLSTM outperforms SimVP in terms of MSE (17.7 vs. 23.8), its reliance on an RNN-based structure results in significantly higher computational cost. SwinLSTM exhibits high FLOPs (69.9G) and lower FPS, making it less efficient for large-scale deployment. This highlights the limitations of recurrent structures in video prediction, whereas PredFormer, with its recurrence-free framework, achieves both higher accuracy and superior efficiency.